OpenStudy (anonymous):

find an exact value of cos 165 degrees help :)?

OpenStudy (anonymous):

i think i might get it but not to sure

OpenStudy (anonymous):

I got it! \[\cos(165) = \cos(120 + 45) = \cos(120)\cos(45) - \sin(120)\sin(45) \] \[= \frac{-1}{2} \frac{1}{\sqrt2}-\frac{\sqrt3}{2}\frac{1}{\sqrt2}\] Simply use: \[\cos(A \pm B) = \cos(A)\cos(B) \mp \sin(A)\sin(B)\]

OpenStudy (anonymous):

\[-\sqrt{6}-\sqrt{2} \] over 4?

OpenStudy (phi):

you could also use the 1/2 angle formula cos(165) = -cos(15) = - cos(30/2)

OpenStudy (anonymous):

i just don't get how to plug in to all this stuff you guys are giving me

OpenStudy (phi):

but your value looks good. you can use a calculator and compare "decimal values" as a check

OpenStudy (anonymous):

when my answer is on a form like i showed above the decimal i get is like.99 something how can that check it?

OpenStudy (phi):

notice you solved cos(15) in your previous problem. you should be -0.96592586 make sure you are in degree mode (and not radian mode)

OpenStudy (anonymous):

and i get .9659 for cos 15 but my answers look like this that i can pick from quantity square root of six plus square root of two divided by four. quantity negative square root of two plus square root of six divided by four. quantity negative square root of six minus square root of two divided by four. quantity negative square root of six plus one divided by four.

OpenStudy (phi):

your answer up above is correct -(sqr(6)+sqr(2))/4 if you use your calculator, this expression will match -0.96592586

OpenStudy (jdoe0001):

in short, do not use the calculator, add up the fractions only, leave it as rational

OpenStudy (anonymous):

ok sweet thanks guys ok now heres were im confused i k now how to fin an exact value for cosine of nineteen pi divided by twelve.

OpenStudy (phi):

It looks like they want you to use sums and differences formulas cos(A+B)= cos(A)cos(B)- sin(A)sin(B) and cos(A-B) = cos(A) cos(B)+ sin(A)sin(B)

OpenStudy (phi):

I would change 19 pi/12 radians into degrees (I think better with degrees)

OpenStudy (anonymous):

how do i know what to plug into the formula that you gave me?

OpenStudy (phi):

First, find 19 pi/12 is 285 degees if you plot this, you see it is 15 degrees past 270